OpenStudy (anonymous):

Wally knows that in order to add or subtract rational expressions, he has to find the least common denominator first. Unfortunately, he cannot remember how to do that. Using complete sentences, explain to Wally how to find least common denominators. Make sure you clearly explain any important items to consider.

OpenStudy (anonymous):

let the denominators be x and y let x=a*b^2*c^3*d, where a,b,c,d,are prime divisors of x let y=a*c^2*d^2*e,where a,c,d,e are prime divisors of y then l.c.d.=a*b^2*c^3*d^2*e